# Scalable Formal Verification via Autoencoder Latent Space Abstraction

### v2

@@ 4.3.2. States and Labels
To ensure that the labeling for each region $r_{i}\in R$ is done correctly in the latent space, we follow the method outlined in (cauchi2019efficiency) and begin by representing $r_{i}$ and its complement $r_{n,i}=X\setminus r_{i}$ with new atomic propositions $p_{i},p_{n,i}$, where $p_{n,i}$ then represents $\neg p_{i}$. The atomic propositions for the NTS are then defined as
@@
For a point $x\in X$, let $q_{x}\ni h_{enc}(x)$ denote the latent region in the partition $Q$ that contains $h_{enc}(x)$, and with an abuse of notation, let $L:X\to 2^{AP}$ or $L:Q\to 2^{AP}$, where APAP is defined in Eq. (8). Under the construction of the latent space described in Section 4.3.1 and labeling procedure described in Section 4.3.2, with high-confidence, it holds that $L(q_{x})\subseteq L(x)$.
Hence we capture (conservatively) all of the regions $R$ (and their negations) by the labels in $\mathcal{N}$, a detailed proof is available in Appendix A.3. We first translate $\varphi$ to NNF as $\mathit{NNF}(\varphi)$ (see Appendix A.1 for details), and then replace every $\neg p_{i}$ in $\mathit{NNF}(\varphi)$ with $p_{n,i}$. Then any trace that satisfies $\overline{\varphi}$ over APAP will also satisfy $\varphi$ over $\Pi$ (cauchi2019efficiency).
@@
$4.$ NTS Transitions: Modeling the Latent Inclusion Dynamics
We propose a novel approach to model inclusion dynamics accurately based on GPR. From an GP model, we then build off of existing methods, e.g., (jackson2021formal; reed2023promises), to generate a NTS abstraction transitions.
@@ 4.4.1. Inclusion GP
Since for each choice of $c\in[a,b]$, $g_{c}$ is a continuous function and each realization of $g(z)$ returns a single element, GPR can be used to learn the $j$-th component of $g_{c}$. However, in our setting we do not have data as $([z_{i},c_{i}],z_{i}^{+})$, but instead have data as $(z_{i},z_{i}^{+})$. To enable GPR, we must learn the unobserved variable $c_{i}$ that matches the data $(z_{i},z_{i}^{+})$. We draw inspiration from Gaussian Process Latent Variable Models (GPLVM) (titsias2010bayesian), which are employed to identify an input data set for GPR when only given an observation dataset. GPLVMs have commonly been used for generative modeling and dimensionality reduction methods, but unlike our approach GPLVMs do not typically have an interpretable latent space which makes predicting the evolution of dynamics infeasible.
Following a GPLVM approach, we first regress over the encoded data $\{(z_{i},z_{i}^{+})\}=\{(h_{enc}(x_{i}),h_{enc}(x_{i}^{+}))\}$ and learn a $c_{i}\in[a,b]$ for each $(z_{i},z_{i}^{+})$ resulting in an augmented dataset $\{([z_{i},c_{i}],z_{i}^{+})\}$. The resulting GP kernel is then $\kappa:\mathbb{R}^{n_{p}+1}\times\mathbb{R}^{n_{p}+1}\rightarrow\mathbb{R}_{\geq 0}$. Using methods outlined in (reed2023promises), we learn a deep kernel model from the augmented data set where the $j$-th component of $g_{c}$ is modeled with $\mu^{(j)}_{D}(z,c)$.
@@
Figure 3. Visualization of Inclusion GP predictions. The initial latent variables $z^{*}$ is shown as a black dot (top-right), observed evolutions $z_{i}^{+}$ as red stars (multiple samples of $h_{enc}(\mathcal{X}^{\prime}_{z^{*}})$), the predicted evolution for 10 $c^{*}$ values of the Inclusion GP with blue rectangles ($\mu^{(j)}_{D}(z^{*},c^{*})\pm\sigma_{D}^{(j)}(z^{*},c^{*})\sqrt{B^{2}-d^{*}}$). :
@@ 4.4.2. NTS Transition Function
For a region $q$ and subset $u$, let $\epsilon^{(j)}_{q,u}=\sup_{z\in q,c\in u}\epsilon^{(j)}(z,c),$ and let the worst case regression error over a region $q$ be $e^{(j)}(q,u)=\sup_{z\in q,c\in u}|\mu_{D}^{(j)}(z,c)-g_{c}^{(j)}(z)|.$ Then, $e^{(j)}(q,u)\leq\epsilon^{(j)}_{q,u}$. The transition from $z\in q$ to $q^{\prime}$ under dynamics $g$ can be over-approximated according to the following proposition, a modified form of (jackson2021formal, Theorem 1) that accounts for the noise-free regression setting.
Proposition 0 ((jackson2021formal), Theorem 1).
@@
Let $\mathcal{R}\subseteq\mathbb{R}^{n_{x}}\times Q$ define a relation such that $(x,q)\in\mathcal{R}$ if $h_{enc}(x)\in q$. NTS $\mathcal{N}$ is a sound abstraction of System (1) in domain $X$, if for any trajectory $\omega_{x_{0}}$, $\exists\omega_{q_{0}}\in\Omega_{q_{0}}$ s.t. $(\omega_{x_{0}}(k),\omega_{q_{0}}(k))\in\mathcal{R}$ and $L(\omega_{q_{0}}(k))\subseteq L(\omega_{x_{0}}(k))$ for all $k\geq 0$.
@@ 4.4.3. RKHS Constants in the Latent Space
As stated in (reed2024error), $d^{*}$ (the parameter used in the GPR error bounded in Proposition 1 and Theorem 8) can be conservatively set to 0 to avoid difficulty in calculating a nontrivial value, as in our case $d^{*}\leq g_{c}^{(j)}(Z)^{T}(K_{(Z,c),(Z,c)}+\sigma_{n}^{2}I)^{-1}g_{c}^{(j)}(Z)$ would require being able to bound the effect of $g_{c}(Z)$ on the quadratic form. $B$ can be formally bounded using approaches in (jackson2021formal), and only depends on having a Lipschitz constant for the dynamics. The Lipschitz constant for $g$ can be easily bounded using the constants for $h_{enc}$ and the dynamics in the original space, or bounded using approaches outlined in (wood1996estimation).
@@ 5. Verification and Refinement
Given that the NTS is a sound abstraction of System (1) per Theorem 11 in the latent space, we can make use of existing methods for verification against LTL specifications to identify the set of states that are guaranteed to satisfy the specification (baier2008principles). Due to the uncertainty introduced by non-determinism and discretization in the latent space, we obtain three sets from verification $Q_{yes}$ (set of states guaranteed to satisfy $\varphi$), $Q_{no}$ (set of states that satisfy $\neg\varphi$), and $Q_{?}=Q\setminus(Q_{yes}\cup Q_{no})$. This uncertainty can result in large volumes of space falling in $Q_{?}$. To address this, we consider iterative refinement of the abstraction and choose states with the most immediate uncertainty to further discretize, i.e., we choose the states in $Q_{?}$ that have transitions to either $Q_{yes}$ or $Q_{no}$ to refine. However, we cannot conclude that states $x$ s.t. $h_{enc}(x)\in Q_{no}$ violate $\varphi$, but states in $Q_{yes}$ are guaranteed to satisfy $\varphi$.
@@ 5.1. Decoding
Then we can use Eq. (16) to identify $\text{Pre}_{h_{enc}}(q)\in X$. The properties of our encoder enforce that the pre-image of a connected set is also connected, hence we can easily over-approximate the pre-image from samples using the same approach described in Section 4.3.1. Alternatively, we can make use of robust pre-image calculation methods for neural networks such as PREMAP (zhang2025premap), which provide a tight under-approximation of the pre-image as a union of polytopes.
@@ 6. Case Studies
In each case study, we first learn $h_{enc}$ from a dataset $D_{enc}$ (the Learning Set from Figure 1) and then optimize the unobserved latent variable for the GP model with a separate dataset $D_{GP}$ (the Regression Set from Figure 1). After generating the latent dataset $\{([z_{i},c_{i}],z_{i}^{+})\}_{i=1}^{m}$, we then learn a deep kernel model that predicts the evolution of the dynamics in the latent space with a network $\psi:\mathbb{R}^{p+1}\rightarrow\mathbb{R}^{p}$ and optimize the base kernel parameters for each dimension as $\kappa(\psi^{(j)}(z,c),\psi^{(j)}(z^{\prime},c^{\prime}))$, following methods outlined in (reed2023promises). The first three examples use the LTL specification $\varphi=\lnot p_{\text{unsafe}}\mathcal{U}p_{\text{goal}}$, where $p_{\text{unsafe}}$ is an atomic proposition that represents unsafe sets including $\mathbb{R}^{n_{x}}\setminus X$, and $p_{\text{goal}}$ is a proposition for the goal. The final example uses a more complex specification of $\varphi=\square(\lnot p_{\text{unsafe}})\wedge\Diamond(p_{\text{goalA}})\wedge\Diamond(p_{\text{goalB}})$ where $p_{\text{goalA}}$ and $p_{\text{goalB}}$ are propositions for two different goal regions.
@@
Figure 4. Results for “3D-Nonlinear-to-2D” Case Study. (a) The 2D latent space with the goal shown outlined in green, the domain outlined in black. (b) The set of states in the latent abstraction that can be guaranteed to reach the goal with the initial abstraction. The result verifies that all latent states reach the goal region and remain in the domain. :
@@
Figure 5. Verification results for the “6D-to-2D” case study. (a) the set of states in the latent abstraction that can be guaranteed to reach the goal with the initial abstraction, (b) after one refinement of the abstraction, and (c) with an obstacle. (c) We plot several trajectories (starting at circles and ending at stars) to verify result, green from $Q_{yes}$, blue from $Q_{?}$, and red from $Q_{no}$. When initializing in $Q_{no}$ we find roughly $40\%$ of trajectories collide with the obstacle and the remaining trajectories are safe and reach the goal. This conservatism comes from over-approximating the unsafe set in the latent space. :
@@
We use 10000 data points to learn the encoder and the DKL model, and 1000 data points for prediction. The encoder is a five layer network with size $\mathbb{R}^{n_{x}}\rightarrow 512\rightarrow 256\rightarrow 128\rightarrow 64\rightarrow\mathbb{R}^{n_{p}}$ with a skip layer adding at the $256$ layer. It takes less than 4 minutes to learn the encoder and the latent deep kernel from data, as well as map the regions of interest. We spend roughly 10 minutes to construct the transition function in the latent space and then less then 1 second is spent to verify against the specification.
@@
Figure 6. Results for the “26D-to-2D” case study. (a) The discretization and convex regions in the 2D latent space of the 26D system. (b) The verification results of the latent abstraction after 10 refinements. Multiple trajectories are plotted with green trajectories satisfying $\varphi$ in the original space and red being unsafe. Note that many safe trajectories from the uncertain region exhibit unsafe behavior in the latent space. (c)-(d) The unsafe and safe latent trajectories that begin in the same region $q$ from (b) shown in the original space of the system with LiDAR measurements shown every 5 time steps.:
@@
We theorize that, given a set of obstacles and a goal position, a 2D latent space could be learned that models this system with minimal non-determinism. We use 10000 data points to learn the encoder, 30000 for the DKL model, and 1000 data points for prediction. It takes roughly 8 minutes to learn the encoder, 5 minutes to map the regions of interest (which is limited by the speed of the simulator), and 1 minute to learn the latent deep kernel. We spend roughly 12 minutes to construct the transition function in the latent space for the initial abstraction and one hour on 10 iterative refinements, with the majority of the time being spent on bounding the outputs of the Inclusion GP model. For each refinement less then 1 second is spent to verify the abstraction against the specification.
@@
As the original function is not smooth (discontinuities in the LiDAR observations occur in domains that are non-convex due to obstacle placement), the function technically does not live in the RKHS of the SE kernel. We instead verify a set of continuous functions that capture the discontinuous evolution. This results in added conservatism in the abstraction, which can be seen with the unverifiable regions, the boundary of which corresponds to regions in the original space that have discontinuities in the dynamics. Alternate kernels that allow for discontinuities may improve the results. It is also possible that using local GP models for each region in which the dynamics are continuous would allow for a better abstraction as the system and approaches such as Work (park2022jump) could be used, but local GP regression is not explored in this work. We also stress that since the original system is discontinuous, it is impossible for a traditional neural network architecture to capture the evolution of the dynamics correctly (as common architectures can only model continuous systems), which further demonstrates the strength of our approach.
@@
Figure 7. Results for the “26D-to-2D” two goal case study. (a) The discretization and convex regions in the 2D latent space of the 26D system. (b) The verification results of the latent abstraction after 5 refinements, with $p_{\text{goalA}}$ in dark green and $p_{\text{goalB}}$ in blue. Multiple trajectories are plotted with green trajectories satisfying $\varphi$ in the original space and red not satisfying. Note that in the uncertain region below $p_{\text{goalB}}$ there are both satisfying and non-satisfying trajectories. (c) An example satisfying trajectory that corresponds to the top most satisfying latent trajectory in (b) shown in the original space of the system with LiDAR measurements shown every 5 time steps, the trajectory is only plotted up to reaching both goals. (d) An non-satisfying trajectory in the original space that corresponds to the red trajectory starting in the top right in (b), the trajectory is plotted for 60 simulated seconds with much of that time spent circling in $p_{\text{goalA}}$.:
@@
We use 10000 data points to learn the encoder, 30000 for the DKL model, and 1000 data points for prediction. It again takes roughly 8 minutes to learn the encoder, 4 minutes to map the regions of interest, and 1 minute to learn the latent deep kernel. After 5 iterative refinements of the latent abstraction, we identify the majority of the space where satisfaction is possible. The results are shown in Figure 7. We note that with a single obstacle, discontinuities in the dynamics from LiDAR measurements have a smaller impact on the abstraction than the previous example, allowing us to identify more of the satisfying space with our conservative, continuous model.

### v3

@@ 4.3.2. States and Labels
To ensure that the labeling for each region $r_{i}\in R$ is done correctly in the latent space, we follow the method outlined in (Cauchi et al., 2019) and begin by representing $r_{i}$ and its complement $r_{n,i}=X\setminus r_{i}$ with new atomic propositions $p_{i},p_{n,i}$, where $p_{n,i}$ then represents $\neg p_{i}$. The atomic propositions for the NTS are then defined as
@@
For a point $x\in X$, let $q_{x}\ni h_{enc}(x)$ denote the latent region in the partition $Q$ that contains $h_{enc}(x)$, and with an abuse of notation, let $L:X\to 2^{AP}$ or $L:Q\to 2^{AP}$, where $AP$ is defined in Eq. (8). Under the construction of the latent space described in Section 4.3.1 and labeling procedure described in Section 4.3.2, then $L(q_{x})\subseteq L(x)$ holds with confidence $\geq(1-\delta)^{|AP|-|L(x)|}$.
Hence we capture (conservatively) all of the regions $R$ (and their negations) by the labels in $\mathcal{N}$, a detailed proof is available in Appendix A.3. We first translate $\varphi$ to NNF as $\mathit{NNF}(\varphi)$ (see Appendix A.1 for details), and then replace every $\neg p_{i}$ in $\mathit{NNF}(\varphi)$ with $p_{n,i}$. Then any trace that satisfies $\overline{\varphi}$ over $AP$ will also satisfy $\varphi$ over $\Pi$ (Cauchi et al., 2019).
@@
4.4. NTS Transitions: Modeling the Latent Inclusion Dynamics
We propose a novel approach to model inclusion dynamics accurately based on GPR. From an GP model, we then build off of existing methods, e.g., (Skovbekk et al., 2025; Reed et al., 2023), to generate a NTS abstraction transitions.
@@ 4.4.1. Inclusion GP
Since for each choice of $c\in[a,b]$, $g_{c}$ is a continuous function and each realization of $g(z)$ returns a single element, GPR can be used to learn the $j$-th component of $g_{c}$. However, in our setting we do not have data as $([z_{i},c_{i}],z_{i}^{+})$, but instead have data as $(z_{i},z_{i}^{+})$. To enable GPR, we must learn the unobserved variable $c_{i}$ that matches the data $(z_{i},z_{i}^{+})$. We draw inspiration from Gaussian Process Latent Variable Models (GPLVM) (Titsias and Lawrence, 2010), which are employed to identify an input data set for GPR when only given an observation dataset. GPLVMs have commonly been used for generative modeling and dimensionality reduction methods, but unlike our approach GPLVMs do not typically have an interpretable latent space which makes predicting the evolution of dynamics infeasible.
Following a GPLVM approach, we first regress over the encoded data $\{(z_{i},z_{i}^{+})\}=\{(h_{enc}(x_{i}),h_{enc}(x_{i}^{+}))\}$ and learn a $c_{i}\in[a,b]$ for each $(z_{i},z_{i}^{+})$ resulting in an augmented dataset $\{([z_{i},c_{i}],z_{i}^{+})\}$. The resulting GP kernel is then $\kappa:\mathbb{R}^{n_{p}+1}\times\mathbb{R}^{n_{p}+1}\rightarrow\mathbb{R}_{\geq 0}$. Using methods outlined in (Reed et al., 2023), we learn a deep kernel model from the augmented data set where the $j$-th component of $g_{c}$ is modeled with $\mu^{(j)}_{D}(z,c)$.
@@
Figure 3. Visualization of Inclusion GP predictions. The initial latent variables $z^{*}$ is shown as a black dot (top-right), observed evolutions $z_{i}^{+}$ as red stars (multiple samples of $h_{enc}(\mathcal{X}^{\prime}_{z^{*}})$), the predicted evolution for 10 $c^{*}$ values of the Inclusion GP with blue rectangles ($\mu^{(j)}_{D}(z^{*},c^{*})\pm\sigma_{D}^{(j)}(z^{*},c^{*})\sqrt{B^{2}-d^{*}}$).
@@ 4.4.2. NTS Transition Function
For a region $q$ and subset $u$, let $\epsilon^{(j)}_{q,u}=\sup_{z\in q,c\in u}\epsilon^{(j)}(z,c),$ and let the worst case regression error over a region $q$ be $e^{(j)}(q,u)=\sup_{z\in q,c\in u}|\mu_{D}^{(j)}(z,c)-g_{c}^{(j)}(z)|.$ Then, $e^{(j)}(q,u)\leq\epsilon^{(j)}_{q,u}$. The transition from $z\in q$ to $q^{\prime}$ under dynamics $g$ can be over-approximated according to the following proposition, a modified form of (Skovbekk et al., 2025, Theorem 1) that accounts for the noise-free regression setting.
Proposition 0 ((Skovbekk et al., 2025), Theorem 1).
@@
Let $\mathcal{R}\subseteq\mathbb{R}^{n_{x}}\times Q$ define a relation such that $(x,q)\in\mathcal{R}$ if $h_{enc}(x)\in q$. NTS $\mathcal{N}$ is a sound abstraction of System (1) in domain $X$, if for any trajectory $\omega_{x_{0}}$, $\exists\omega_{q_{0}}\in\Omega_{q_{0}}$ s.t. $(\omega_{x_{0}}(k),\omega_{q_{0}}(k))\in\mathcal{R}$ and $L(\omega_{q_{0}}(k))\subseteq L(\omega_{x_{0}}(k))$ holds with confidence at least $(1-\delta)^{\max_{x}|AP|-|L(x)|}$ for all $k\geq 0$.
@@ 4.4.3. RKHS Constants in the Latent Space
As stated in (Reed et al., 2025), $d^{*}$ (the parameter used in the GPR error bounded in Proposition 1 and Theorem 8) can be conservatively set to 0 to avoid difficulty in calculating a nontrivial value, as in our case $d^{*}\leq g_{c}^{(j)}(Z)^{T}(K_{(Z,c),(Z,c)}+\sigma_{n}^{2}I)^{-1}g_{c}^{(j)}(Z)$ would require being able to bound the effect of $g_{c}(Z)$ on the quadratic form. $B$ can be formally bounded using approaches in (Skovbekk et al., 2025), and only depends on having a Lipschitz constant for the dynamics. The Lipschitz constant for $g$ can be easily bounded using the constants for $h_{enc}$ and the dynamics in the original space, or bounded using approaches outlined in (Wood and Zhang, 1996).
@@ 5. Verification and Refinement
Given that the NTS is a sound abstraction of System (1) per Theorem 11 in the latent space, we can make use of existing methods for verification against LTL specifications to identify the set of states that are guaranteed to satisfy the specification (Baier and Katoen, 2008). Due to the uncertainty introduced by non-determinism and discretization in the latent space, we obtain three sets from verification $Q_{yes}$ (set of states guaranteed to satisfy $\varphi$), $Q_{no}$ (set of states that satisfy $\neg\varphi$), and $Q_{?}=Q\setminus(Q_{yes}\cup Q_{no})$. This uncertainty can result in large volumes of space falling in $Q_{?}$. To address this, we consider iterative refinement of the abstraction and choose states with the most immediate uncertainty to further discretize, i.e., we choose the states in $Q_{?}$ that have transitions to either $Q_{yes}$ or $Q_{no}$ to refine. However, we cannot conclude that states $x$ s.t. $h_{enc}(x)\in Q_{no}$ violate $\varphi$, but states in $Q_{yes}$ are guaranteed to satisfy $\varphi$.
@@ 5.1. Decoding
Then we can use Eq. (16) to identify $\text{Pre}_{h_{enc}}(q)\in X$. The properties of our encoder enforce that the pre-image of a connected set is also connected, hence we can easily over-approximate the pre-image from samples using the same approach described in Section 4.3.1. Alternatively, we can make use of robust pre-image calculation methods for neural networks such as PREMAP (Zhang et al., 2025), which provide a tight under-approximation of the pre-image as a union of polytopes.
@@ 6. Case Studies
In each case study, we first learn $h_{enc}$ from a dataset $D_{enc}$ (the Learning Set from Figure 1) and then optimize the unobserved latent variable for the GP model with a separate dataset $D_{GP}$ (the Regression Set from Figure 1). After generating the latent dataset $\{([z_{i},c_{i}],z_{i}^{+})\}_{i=1}^{m}$, we then learn a deep kernel model that predicts the evolution of the dynamics in the latent space with a network $\psi:\mathbb{R}^{p+1}\rightarrow\mathbb{R}^{p}$ and optimize the base kernel parameters for each dimension as $\kappa(\psi^{(j)}(z,c),\psi^{(j)}(z^{\prime},c^{\prime}))$, following methods outlined in (Reed et al., 2023). The first three examples use the LTL specification $\varphi=\lnot p_{\text{unsafe}}\mathcal{U}p_{\text{goal}}$, where $p_{\text{unsafe}}$ is an atomic proposition that represents unsafe sets including $\mathbb{R}^{n_{x}}\setminus X$, and $p_{\text{goal}}$ is a proposition for the goal. The final example uses a more complex specification of $\varphi=\square(\lnot p_{\text{unsafe}})\wedge\Diamond(p_{\text{goalA}})\wedge\Diamond(p_{\text{goalB}})$ where $p_{\text{goalA}}$ and $p_{\text{goalB}}$ are propositions for two different goal regions. When over-approximating a region, we use 200000 samples and set $\delta=0.01$ for a confidence of 99%.
@@
Figure 4. Results for “3D-Nonlinear-to-2D” Case Study. (a) The 2D latent space with the goal shown outlined in green, the domain outlined in black. (b) The set of states in the latent abstraction that can be guaranteed to reach the goal with the initial abstraction. The result verifies that all latent states reach the goal region and remain in the domain.
A partition of the 2D latent space, shown with a multi colored grid, with the goal region outlined in green and the domain boundary in black. The next image shows that the entire latent space is verified to satisfy reaching the goal and remaining in the domain.
@@
Figure 5. Verification results for the “6D-to-2D” case study. (a) the set of states in the latent abstraction that can be guaranteed to reach the goal with the initial abstraction, (b) after one refinement of the abstraction, and (c) with an obstacle. (c) We plot several trajectories (starting at circles and ending at stars) to verify result, green from $Q_{yes}$, blue from $Q_{?}$, and red from $Q_{no}$. When initializing in $Q_{no}$ we find roughly $40\%$ of trajectories collide with the obstacle and the remaining trajectories are safe and reach the goal. This conservatism comes from over-approximating the unsafe set in the latent space.
Verification results for the 6D nonlinear example shown in the 2D latent space. the first two figures show that the entire latent domain can be verified as reaching the target set while remaining in the domain. The final figure shows results with an obstacle in the domain, with trajectories from the verified satisfying region are shown reaching the goal, the uncertain region also reaches the goal. The unsatisfying regions are shown both colliding with the obstacle or reaching the goal.
@@
We use 10000 data points to learn the encoder and the DKL model, and 1000 data points for prediction. The encoder is a five layer network with size $\mathbb{R}^{n_{x}}\rightarrow 512\rightarrow 256\rightarrow 128\rightarrow 64\rightarrow\mathbb{R}^{n_{p}}$ with a skip layer adding at the $256$ layer. It takes 187.64 seconds to learn the encoder and the latent deep kernel from data, as well as map the regions of interest. We spend 564.26 seconds to construct the transition function in the latent space and then less then 1 second is spent to verify against the specification.
@@
Figure 6. Results for the “26D-to-2D” case study. (a) The discretization and convex regions in the 2D latent space of the 26D system. (b) The verification results of the latent abstraction after 10 refinements. Multiple trajectories are plotted with green trajectories satisfying $\varphi$ in the original space and red being unsafe. Note that many safe trajectories from the uncertain region exhibit unsafe behavior in the latent space. (c)-(d) The unsafe and safe latent trajectories that begin in the same region $q$ from (b) shown in the original space of the system with LiDAR measurements shown every 5 time steps.
Verification in a 2D latent space of a 26D system, the abstraction correctly represents uncertainty due to inclusion dynamics and over-approximated unsafe sets
@@
We theorize that, given a set of obstacles and a goal position, a 2D latent space could be learned that models this system with minimal non-determinism. We use 10000 data points to learn the encoder, 30000 for the DKL model, and 1000 data points for prediction. It takes 515.298 seconds to learn the encoder and 67.47 seconds to learn the latent deep kernel. We spend 681.23 seconds to construct the initial latent abstraction and two hours on 10 iterative refinements, with the majority of the time being spent on bounding the outputs of the Inclusion GP model. For each refinement less then 1 second is spent to verify the abstraction against the specification.
@@
As the original function is not smooth (discontinuities in the LiDAR observations occur in domains that are non-convex due to obstacle placement), the function technically does not live in the RKHS of the SE kernel. We instead verify a set of continuous functions that capture the discontinuous evolution. This results in added conservatism in the abstraction, which can be seen with the unverifiable regions, the boundary of which corresponds to regions in the original space that have discontinuities in the dynamics. Alternate kernels that allow for discontinuities may improve the results. It is also possible that using local GP models for each region in which the dynamics are continuous would allow for a better abstraction as the system and approaches such as Work (Park, 2022) could be used, but local GP regression is not explored in this work. We also stress that since the original system is discontinuous, it is impossible for a traditional neural network architecture to capture the evolution of the dynamics correctly (as common architectures can only model continuous systems), which further demonstrates the strength of our approach.
@@
Figure 7. Results for the “26D-to-2D” two goal case study. (a) The discretization and convex regions in the 2D latent space of the 26D system. (b) The verification results of the latent abstraction after 5 refinements, with $p_{\text{goalA}}$ in dark green and $p_{\text{goalB}}$ in blue. Multiple trajectories are plotted with green trajectories satisfying $\varphi$ in the original space and red not satisfying. Note that in the uncertain region below $p_{\text{goalB}}$ there are both satisfying and non-satisfying trajectories. (c) An example satisfying trajectory that corresponds to the top most satisfying latent trajectory in (b) shown in the original space of the system with LiDAR measurements shown every 5 time steps, the trajectory is only plotted up to reaching both goals. (d) An non-satisfying trajectory in the original space that corresponds to the red trajectory starting in the top right in (b), the trajectory is plotted for 60 simulated seconds with much of that time spent circling in $p_{\text{goalA}}$.
Verification in a 2D latent space of a 26D system, the abstraction correctly identifies regions that satisfy the specification.
@@
We use 10000 data points to learn the encoder, 30000 for the DKL model, and 1000 data points for prediction. After 5 iterative refinements of the latent abstraction, we identify the majority of the space where satisfaction is possible. The results are shown in Figure 7. We note that with a single obstacle, discontinuities in the dynamics from LiDAR measurements have a smaller impact on the abstraction than the previous example, allowing us to identify more of the satisfying space with our conservative, continuous model.